OpenStudy (watermelon14):

Check: is the vertex form of -3x^2-6x+3.....this:(x+1)^2+6

OpenStudy (watermelon14):

@radar

OpenStudy (watermelon14):

@zepdrix

zepdrix (zepdrix):

Hmm close :d

zepdrix (zepdrix):

Your vertex form should look like this:\[\large\rm y=\color{orangered}{a}(x-h)^2+k\]Where your vertex is located at (h,k). Yes, you have the correct vertex, but you forgot about your \(\large\rm \color{orangered}{a}\).

OpenStudy (watermelon14):

yeah, how do i find a

zepdrix (zepdrix):

\[\Large\rm \color{blue}{a}x^2+\color{red}{b}x+\color{green}{c}=\color{blue}{-3}x^2+\color{red}{-6}x+\color{green}{3}\]

zepdrix (zepdrix):

Oh I should have colored the a in orange again, not blue :) lol

zepdrix (zepdrix):

Anyway, it's the same a that you started with.

zepdrix (zepdrix):

The same a that you used to find the x-coordinate of your vertex in the -(b/2a) formula

OpenStudy (watermelon14):

-3?

zepdrix (zepdrix):

yup! :)\[\large\rm y=\color{orangered}{-3}(x+1)^2+6\]

OpenStudy (watermelon14):

yasss. one more question how do you find the roots to this problem

zepdrix (zepdrix):

Oh I was waiting for you to post a problem. haha, roots to the same problem? XD

OpenStudy (watermelon14):

haha do you want me to open another question

zepdrix (zepdrix):

`Roots` is another fancy word that we use. Remember earlier I was calling them `zeros` or `x-intercepts`. All three of those terms mean the same thing. So recall, to find roots/zeros/solutions, we let y=0.

zepdrix (zepdrix):

\[\large\rm 0=-3(x+1)^2+6\]And solve for x.

zepdrix (zepdrix):

Understand how to do that? :O

OpenStudy (watermelon14):

i think , i am trying to work it out right now

OpenStudy (watermelon14):

wait can't i just set -3x^2-6x+3

OpenStudy (watermelon14):

equal to zero

zepdrix (zepdrix):

Yes, and use the `quadratic formula` from there. That is your other option. :)

zepdrix (zepdrix):

It's a little bit easier to solve the vertex form than it is to simplify the quadratic formula. Both are quite fine though. Whichever you're more familiar with.

OpenStudy (watermelon14):

fabulous, i got something weird like \[1\pm \sqrt{17}/-1\]

zepdrix (zepdrix):

Hmm that square root doesn't look quite right +_+

zepdrix (zepdrix):

Which method did you use?

OpenStudy (watermelon14):

i did the quadratic formula of -3x^2-6x+3

zepdrix (zepdrix):

\[\large\rm x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Then plugging in the numbers,\[\large\rm x=\frac{6\pm\sqrt{(-6)^2-4(-3)(3)}}{2(-3)}\]it should look like that, ya?

OpenStudy (watermelon14):

i'll do the rest

OpenStudy (watermelon14):

\[\left( 6\pm \sqrt{102} \right)/-6\]

zepdrix (zepdrix):

\[\large\rm x=\frac{6\pm\sqrt{(-6)^2-4(-3)(3)}}{2(-3)}\] \[\large\rm x=\frac{6\pm\sqrt{36+36}}{-6}\] \[\large\rm x=\frac{6\pm\sqrt{2\cdot36}}{-6}\]Hmm 106? :o

OpenStudy (watermelon14):

omg no 72

zepdrix (zepdrix):

We would like to simplify the root, so it's actually preferable to write it this way :)\[\large\rm \sqrt{36+36}\quad=\quad\sqrt{2\cdot36}\quad=\quad\sqrt{2}\cdot\sqrt{36}\]

zepdrix (zepdrix):

\[\large\rm x=\frac{6\pm\sqrt{2}\cdot\sqrt{36}}{-6}\]Do you see how it will simplify further?

OpenStudy (watermelon14):

yes eventually it will be \[1\pm \sqrt{12}/-1\] right?

zepdrix (zepdrix):

12? :o hmm

OpenStudy (watermelon14):

i divide 72 by 6

zepdrix (zepdrix):

And if you're going to write it like that, make sure you place brackets around your numerator,\[(1\pm \sqrt{12})/-1\]But no not 12 :(

zepdrix (zepdrix):

Divide 72 by 36, not 6.

zepdrix (zepdrix):

If you divide it by 6,\[\large\rm \sqrt{72}\quad=\quad \sqrt{12\cdot6}\]Then there is no way to get the 6 out of the square root.

OpenStudy (watermelon14):

2

zepdrix (zepdrix):

\[\large\rm x=\frac{6\pm\sqrt{2}\cdot\sqrt{36}}{-6}\] \[\large\rm x=\frac{6\pm\sqrt{2}\cdot6}{-6}\]Dividing everything by 6,\[\large\rm x=\frac{1\pm\sqrt{2}}{-1}\quad=\quad -1\pm\sqrt{2}\]Ah yes, 2 looks better! :)

zepdrix (zepdrix):

That was a lot of nasty algebra to get there. Let's remember what the question was asking for. Roots, yes? So we have two of them: One corresponding to the positive sqrt(2) and the other for the negative sqrt(2).\[\large\rm x=-1+\sqrt2\qquad\qquad and\qquad\qquad x=-1-\sqrt2\]

OpenStudy (watermelon14):

fab, so we are done?

zepdrix (zepdrix):

yes :)

zepdrix (zepdrix):

now go eat a strawberry :U even a blueberry will do. just not the watermelon, i beg of you!! >.< lol

OpenStudy (watermelon14):

hahaha fine ill eat watermelon tomorrow :) thanks for all the help!!!